Consider a firm with a contract to sell an asset for $220,000 seven years from now. The asset costs $75,000 to produce today. Given a relevant discount rate on this asset of 7.5 percent per year, will the firm make a profit on this asset. At what rate does the firm just break even?
Answer #1

Hello SIr/ Mam

Given that:

Cost of producing the asset today = $75,000

Contract of selling price = $220,000

Discount Rate = 7.5%

n = 7

FV = $75,000* (1 +7.5%) = $1,24, 428.69

Hence,

Profit = $220,000 - $1, 24, 428.69 = $95,571.31

Now, the breakeven rate =

Using excel function, "=RATE(7,0,-75000,220000,0)", we get Rate = 16.62%

I hope this solves your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.
Answer #2

FV of $75,000

 FV= 75000*(1+7.5%)^7=$124,428.69

Profit

 220,000-124,428.69= $95,571.31

=> Firm will make a profit on this asset

The rate dose the firm just break even:

FV=PV(1+r)^t

220,000=75,000*(1+r)^7

=> r= 16.62%=> This is the rate at which the firm just break even.

 


answered by: trang trần